Find a power series for the function, centered at c, and determine the interval of convergence.x +1 4x – 7 c) f(x)= 2x2 + 3x – 2 ; c=0 d) -; 2x² + 5x – 3
Answer #1

c) 1 f(a) = 40-7 222 +30-2 f (x) = 4a-u7 (227) (4+2) fiad ) · 2 2a-l 1+2 f(x)= 2 (1-22) 3 211+ 32) f(a)= a (1-25 + 3 (1+%25 2f(x) = + 1/7 +2 (x-2) 1 + 213 2/7 5 (1 + (2-2) PCIE n + find = ! (3 (42) 2 35 7 For internal of convergence اے 3 3 (**)) < |<